ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 62 Hua Liang, Chinese Academy of Sciences, China and Humboldt-Universität zu Berlin, Germany Partially linear errors-in-variables model We consider the partially linear model when the linear predictors are measured with additive error. A simple modification of the likelihhod estimate is derived. The resulting estimator is shown to be consistent and its asymptotic distribution theory is derived. Consistent standard error estimates using sandwich–type ideas are also developed. References: R. -J. Carroll, D. Ruppert and L. -A. Stefanski, Nonlinear Measurement Error Models. Chapman and Hall, New York, 1995.